OpenStudy (anonymous):

@phi Is the system consistent or inconsistent? 2x + 3y = 5 and 4x + 6y = 19

OpenStudy (anonymous):

@phi Please Help!

OpenStudy (phi):

can you change both equations to y= mx + b form?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

just a sec

OpenStudy (phi):

the other way to do this problem is divide the second equation by 2 (that means divide both sides and *all terms* by 2) you start with 4x + 6y = 19 now divide both sides and all terms by 2 2x + 3y = 19/2 compare to the first equation 2x + 3y = 5 the left sides are the same , but the numbers on the right side are different. common sense says: if 19/2 = 2x+3y and 2x+3y= 5 then 19/2 = 5 which cannot be true. so this system is inconsistent

OpenStudy (anonymous):

Than You, I might Have A few More Questions

OpenStudy (phi):

the other way is to change to y= mx + b form for both lines. you will find they have the *same slope* - that means either they are parallel and never meet or are the same line. The "b" will be different. that means different lines, and you have option 1: the lines are parallel and never meet. so no solution (i.e. inconsistent)

OpenStudy (anonymous):

Thank you, I understand now, but I have a new Question posted :)